Let T: R^3→R^2 be given by:

T(x2,x2,x3)= (X1+x2+x3,x2+x3)

Prove that T is a linear transformation. Also find the rank and nullity of T.


1
Expert's answer
2021-02-07T15:11:57-0500

T: R3→R2 be given by:

T(x1,x2,x3)=(x1+x2+x3,x2+x3)

"\implies" First, we need to check T is a linear transformation.

let (x1,x2,x3), (y1,y2,y3) "\in" R3 and "\alpha" "\isin" R

T("\alpha" (x1,x2,x3) + (y1,y2,y3)) = T(("\alpha"x1,"\alpha"x2,"\alpha"x3) + (y1,y2,y3))

= T( "\alpha"x1+y1, "\alpha"x2+y2, "\alpha"x3+y3)

= ("\alpha"x1+y1+"\alpha"x2+y2+"\alpha"x3+y3 , "\alpha"x2+y2+"\alpha"x3+y3)

= ("\alpha"(x1+x2+x3) + (y1+y2+y3), "\alpha"(x2+x3)+(y2+y3))

= "\alpha"(x1+x2+x3,x2+x3) + (y1+y2+y3,y2+y3)

= "\alpha"T(x1,x2,x3) + T(y1,y2,y3)

since,it is satisfying the definition of linear transformation.

Hence,

T is a linear transformation.


"\implies" Now we need to find the rank and nullity of T.

as given;

T(x1,x2,x3)=(x1+x2+x3,x2+x3)

here first to find the matrix corresponding to this linear transformation,

let "\beta" ={(1,0,0),(0,1,0),(0,0,1)}

T(1,0,0) = (1,0) = 1(1,0) + 0(0,1)

T(0,1,0) = (1,1) = 1(1,0) + 1(0,1)

T(0,0,1) = (1,1) = 1(1,0) + 1(0,1)


[T] ="\begin{bmatrix}
 1 & 1 & 1 \\
 0 & 1 & 1
\end{bmatrix}"


since matrix have two non-zero rows. therefore rank of a matrix is 2.

"\implies" rank (T)= 2

using rank nullity theorem which states that rank(T) + nullity(T) = n;

rank(T) + nulity(T) = 3 "\because" n = 3

nulity(T) = 3 - rank(T) = 3 - 2 =1

"\implies" nulity(T) = 1